PLEASE SHOW WORK:

Assets, Inc. plans to issue $5 million of bonds with a coupon rate of 7 percent, a par value of $1,000, semiannual coupons, and 30 years to maturity. The current market interest rate on these bonds is 6 percent. In one year, the interest rate on the bonds will be either 9 percent or 5 percent with equal probability. Assume investments are risk-neutral.

a.) If the bonds are noncallable, what is the price of the bonds today?

b.) If the bonds are callable one year from today at $1,080, will their price be greater or less than the price you computed in (a)? Why?
Answer #1
a) Price of bond if non -callable
Number of bond issued =5million/1000 5000
Face Value of each bond $1,000
Expected interest rate after one year=0.5*(9+5) 7%
Rate Expected semi annual interest =7/2 3.50%
Nper Number of semi annual periods after 1 year                  58 (29*2)
Pmt Amount of semi annual coupon payment=1000*(7/2)% $35.00
Fv Payment at maturity 1000.00
PV Expected Bond Price after one year $1,000.00
(Using PV function of excel)
Rate Current Semi annual Market Interest=6/2= 3%
Nper Number of Coupon payment in one year                    2
Pmt Amount of semi annual coupon payment=1000*(7/2)% $35.00
Fv Bond Price after one year $1,000.00
PV Current Bond Price = $1,009.57
(Using PV function of excel)
b) Price of bond if callable
Rate Current Semi annual Market Interest=6/2= 3%
Nper Number of Coupon payment in one year                    2
Pmt Amount of semi annual coupon payment=1000*(7/2)% $35.00
Fv Bond Call Price after one year $1,080.00
PV Current Bond Price = $1,084.98
(Using PV function of excel)
Their price will be greater than the price computed in (a)
Because the Call amount is higher than the expected price of bond after one year

F15 - X B Fax C =PV(F11, F12,-F13,-1000) D 2 A F G Nm two como Rate Nper Pmt Price of bond if non-callable Number of bond issIle TIME THISCIL Page LayouLTUUlas DalaNEVIEW VIEW TICIP Tell Me Will you wa IL LU UU F21 x B for C A G 17 18 =PV(F17,F18,-F1529 : * B v fac C =PV(F25,526,-F27-F28) D A F 13 Pmt 14 FV $35.00 1000.00 $1,000.00 PV 3% Rate Nper Pmt Amount of semi annual